# Towards Interpretable Hate Speech Detection using Large Language Model-extracted Rationales

### v1

@@ Abstract
Although social media platforms are a prominent arena for users to engage in interpersonal discussions and express opinions, the facade and anonymity offered by social media may allow users to spew hate speech and offensive content. Given the massive scale of such platforms, there arises a need to automatically identify and flag instances of hate speech. Although several hate speech detection methods exist, most of these black-box methods are not interpretable or explainable by design. To address the lack of interpretability, in this paper, we propose to use state-of-the-art Large Language Models (LLMs) to extract features in the form of rationales from the input text, to train a base hate speech classifier, thereby enabling faithful interpretability by design. Our framework effectively combines the textual understanding capabilities of LLMs and the discriminative power of state-of-the-art hate speech classifiers to make these classifiers faithfully interpretable. Our comprehensive evaluation on a variety of social media hate speech datasets demonstrate: (1) the goodness of the LLM-extracted rationales, and (2) the surprising retention of detector performance even after training to ensure interpretability.
@@
Ayushi $*{}^{*}$    Amrita $*{}^{*}$    Paras Sheth    Huan Liu School of Computing and Augmented Intelligence Arizona State University {anirmal1, abhatt43, psheth5, huanliu}@asu.edu
@@ 1 Introduction
    We evaluate the godness of LLM-extracted features and rationales, and measure the alignment of such with human annotated rationales.
@@ 3.1 Datasets
Table 1: Dataset statistics for explicit hate speech dataset from different social media platforms.
@@ 3.2 Baselines
CATCH Furthermore, we compare our model with CATCH  Sheth et al. (2023a) framework which disentangles the input representations into invariant and platform-dependent features.
@@ 4.1 Performance of ChatGPT on the hate speech detection task
Text
"always thought it astounding no policing agency dares do this to the armed muslim compounds all over the country"
@@
Table 3: Evaluation results for our SHIELD framework vs. the baseline models. Implicit HS refers to the Implicit Hate Speech Corpus. Values in bold denote the best performance, and underlined values denotes the second-best performance.
@@ 4.2 Goodness of ChatGPT extracted features or rationales
We are interested to evaluate the textual and contextual understanding capabilities of ChatGPT in order to extract features in the form of rationales from the input text that are meaningful to the task of hate speech detection. Following a similar construction as in  Jain et al. (2020), we use the LLM (i.e., GPT-3.5) as the extractor model, which unlike the extractor model in  Jain et al. (2020), does not require any additional task-specific fine-tuning. This is possible due to the instruction-following capabilities of recent LLMs. We carefully craft a prompt (as shown in Table 4) to extract cuss words, derogatory language and rationales from the input text that serve as interpretable features that can be used in the subsequent predictor model (HateBERT) in order to have a faithfully interpretable hate speech detector. In order to evaluate the goodness of the extracted features or rationales, we compare ChatGPT-extracted rationales with human-annotated ground truth rationales. We use the annotated rationale spans in the HateXplain Mathew et al. (2021) dataset. After some standard pre-processing such as removing stop words, we compute the similarity between the ChatGPT extracted rationales for the input text from HateXplain dataset and the human-annotated rationales and report these scores in Table 5. We see significant overlap and a high semantic similarity between the LLM and human rationales.
@@
We also present some examples in Figure 2 to qualitatively discern the overlap between the human-annotated rationales and the LLM-extracted ones. Text in red are rationales annotated by human annotators, text in blue are rationales or words identified by the LLM and text in purple are the spans where both the LLM and human annotations overlap. From these exmaples, we see that there is overall a high degree of overlap, and the LLM is able to capture semantically relevant portions of the text. Interestingly, we also see that while human annotators often annotate words or spans with lesser relevance to the task, the LLM extracted rationales do not contain these spans (such as ‘aids figures’ and ‘prominently’ in the first example in Figure 2). Using LLM-extracted rationales for training might be even more useful in such cases since some of the noisy signals in the data can be avoided.
@@ 4.3 Hate speech detector performance after training with extracted rationales
Similarity Metric Similarity Coefficients
Overlap Similarity 52.38%
Cosine Similarity 74.51%
Table 5: Similarity Metric between HateXplain human explanations and LLMs explanations.
@@ 5.2 LLMs as Experts or Feature Extractors
Recent advancements in LLM research have demonstrated improved performance across not only many natural language tasks Min et al. (2023), but also more challenging domains such as writing and debugging code, performing mathematical reasoning Bubeck et al. (2023), etc. This has motivated a line of research where the community has been trying to evaluate how well these LLMs can perform different tasks. LLMs have shown promise in the task of data annotation He et al. (2023); Bansal and Sharma (2023), information extraction  Dunn et al. (2022), and even reasoning  Ho et al. (2022). Given the ease with which these LLMs can be queried, these models often serve as faulty experts or pseudo oracles in many tasks. Past exploration has investigated whether language models can be used as factual knowledge bases Petroni et al. (2019). A recent work has explored the possibility of using LLMs in the hate speech detection task Kumarage et al. (2024). Similar to our approach, authors in  Hasanain et al. (2023) have tried to perform propaganda span annotation using language models. However, our approach focuses on leveraging the extracted spans, words and rationales to augment a detector model to enable interpretability in an otherwise black-box model.
@@ 6 Conclusion and Future Work
While our work follows that of  Jain et al. (2020) and we establish faithfulness by construction, future work could explore better ways to evaluate the faithfulness of the resulting detector. In this work, we verified the goodness of the extracted rationales by comparing it with the ground truth for one dataset. Future work can investigate better automated ways to evaluate and verify the quality of the LLM-extracted rationales.

### v2

@@ Abstract
Although social media platforms are a prominent arena for users to engage in interpersonal discussions and express opinions, the facade and anonymity offered by social media may allow users to spew hate speech and offensive content. Given the massive scale of such platforms, there arises a need to automatically identify and flag instances of hate speech. Although several hate speech detection methods exist, most of these black-box methods are not interpretable or explainable by design. To address the lack of interpretability, in this paper, we propose to use state-of-the-art Large Language Models (LLMs) to extract features in the form of rationales from the input text, to train a base hate speech classifier, thereby enabling faithful interpretability by design. Our framework effectively combines the textual understanding capabilities of LLMs and the discriminative power of state-of-the-art hate speech classifiers to make these classifiers faithfully interpretable. Our comprehensive evaluation on a variety of English language social media hate speech datasets demonstrate: (1) the goodness of the LLM-extracted rationales, and (2) the surprising retention of detector performance even after training to ensure interpretability. All code and data will be made available at https://github.com/AmritaBh/shield.
@@
Ayushi Nirmal    Amrita Bhattacharjee    Paras Sheth    Huan Liu School of Computing and Augmented Intelligence Arizona State University {anirmal1, abhatt43, psheth5, huanliu}@asu.edu
@@ 1 Introduction
    We evaluate the goodness of LLM-extracted features and rationales, and measure the alignment of such with human annotated rationales.
@@ 3.1 Datasets
Implicit HS 20,391 7,100 34.8
Table 1: Dataset statistics for explicit and implicit hate speech datasets comprising data from different social media platforms.
@@ 3.2 Baselines
CATCH: Furthermore, we compare our model with CATCH  Sheth et al. (2023a) framework which disentangles the input representations into invariant and platform-dependent features.
@@ 4.1 Performance of ChatGPT on the hate speech detection task
Text "always thought it astounding no policing agency dares do this to the armed muslim compounds all over the country"
@@
Table 3: Evaluation results (test set accuracy) for our SHIELD framework vs. the baseline models. Implicit HS refers to the Implicit Hate Speech Corpus. Values in bold denote the best performance, and underlined values denotes the second-best performance.
@@ 4.2 Goodness of ChatGPT extracted features or rationales
We are interested to evaluate the textual and contextual understanding capabilities of ChatGPT in order to extract features in the form of rationales from the input text that are meaningful to the task of hate speech detection. Following a similar construction as in  Jain et al. (2020), we use the LLM (i.e., GPT-3.5) as the extractor model, which unlike the extractor model in  Jain et al. (2020), does not require any additional task-specific fine-tuning. This is possible due to the instruction-following capabilities of recent LLMs. We carefully craft a prompt (as shown in Table 4) to extract cuss words, derogatory language and rationales from the input text that serve as interpretable features that can be used in the subsequent predictor model (HateBERT) in order to have a faithfully interpretable hate speech detector. In order to evaluate the goodness of the extracted features or rationales, we compare ChatGPT-extracted rationales with human-annotated ground truth rationales. We use the annotated rationale spans in the HateXplain Mathew et al. (2021) dataset. After some standard pre-processing such as removing stop words, we compute the similarity between the ChatGPT extracted rationales for the input text from HateXplain dataset and the human-annotated rationales and report these scores in Table 5. We compute similarity metrics in both the token space (Jaccard and Overlap similarity) and in the latent space (Cosine and Semantic similarity with Universal Sentence Encoder embeddings Cer et al. (2018)) We see significant overlap and a high semantic similarity between the LLM and human rationales.
@@
We also present some examples in Figure 2 to qualitatively discern the overlap between the human-annotated rationales and the LLM-extracted ones. Text in red are rationales annotated by human annotators, text in blue are rationales or words identified by the LLM and text in purple are the spans where both the LLM and human annotations overlap. From these examples, we see that there is overall a high degree of overlap, and the LLM is able to capture semantically relevant portions of the text. Interestingly, we also see that while human annotators often annotate words or spans with lesser relevance to the task, the LLM extracted rationales do not contain these spans (such as ‘aids figures’ and ‘prominently’ in the first example in Figure 2). Using LLM-extracted rationales for training might be even more useful in such cases since some of the noisy signals in the data can be avoided.
@@ 4.3 Hate speech detector performance after training with extracted rationales
Similarity Metric Similarity Coefficients (%)
Jaccard Similarity 60.39
Overlap Similarity 99.17
Cosine Similarity 74.51
Semantic Similarity (via USE) 56.09
Table 5: Similarity between HateXplain human explanations and LLM-extracted features/rationales.
@@ 5.2 LLMs as Experts or Feature Extractors
Recent advancements in LLM research have demonstrated improved performance across not only many natural language tasks Min et al. (2023), but also more challenging domains such as writing and debugging code, performing mathematical reasoning Bubeck et al. (2023), etc. This has motivated a line of research where the community has been trying to evaluate how well these LLMs can perform different tasks. LLMs have shown promise in the task of data annotation He et al. (2023); Bansal and Sharma (2023), information extraction  Dunn et al. (2022), text classification Kocoń et al. (2023); Bhattacharjee and Liu (2024), and even reasoning  Ho et al. (2022). Given the ease with which these LLMs can be queried, these models often serve as faulty experts or pseudo oracles in many tasks. Past exploration has investigated whether language models can be used as factual knowledge bases Petroni et al. (2019). A recent work has explored the possibility of using LLMs in the hate speech detection task Kumarage et al. (2024). Similar to our approach, authors in  Hasanain et al. (2023) have tried to perform propaganda span annotation using language models. However, our approach focuses on leveraging the extracted spans, words and rationales to augment a detector model to enable interpretability in an otherwise black-box model.
@@ 6 Conclusion and Future Work
While our work follows that of  Jain et al. (2020) and we establish faithfulness by construction, future work could explore better ways to evaluate the faithfulness of the resulting detector. In this work, we verified the goodness of the extracted rationales by comparing it with the ground truth for one dataset. Future work can investigate better automated ways to evaluate and verify the quality of the LLM-extracted rationales. Furthermore, an interesting and responsible direction forward would be the development of hybrid approaches that leverage LLMs for extracting rationales at scale and then employing human experts to verify the validity and quality of these rationales. This would also alleviate some of the concerns surrounding LLM hallucinations and biases in the LLM being propagated into the rationale extraction step.